OpenStudy (anonymous):

The second term in a geometric sequence is 20. The fourth term in the same sequence is 45/4, or 11.25. What is the common ration in this sequence?

jimthompson5910 (jim_thompson5910):

second term: 20 third term: 20r fourth term: (20r)*r = 20r^2 the fourth term is also 11.25, so 20r^2 = 11.25 solve for r

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

I got 0.75.

jimthompson5910 (jim_thompson5910):

r = 0.75 or r = 3/4 good

jimthompson5910 (jim_thompson5910):

now because we have an r^2 term, there's a possibility r could be negative as well so we have to include r = -3/4

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

so in the end r = 3/4 or r = -3/4

jimthompson5910 (jim_thompson5910):

yw